OpenStudy (anonymous):

The cost to produce a product is modeled by the function f(x) = 5x^2 − 70x + 258 where x is the number of products produced. Complete the square to determine the minimum cost of producing this product. 5(x − 7)^2 + 13; The minimum cost to produce the product is $13. 5(x − 7)^2 + 13; The minimum cost to produce the product is $7. 5(x − 7)^2 + 258; The minimum cost to produce the product is $7. 5(x − 7)^2 + 258; The minimum cost to produce the product is $258.

OpenStudy (anonymous):

the third one..

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

ok... the first step: Okay, it would be good to rewrite this in the form: x^2 -14x + 258 / 5 = 0 Do you see how I got to this step?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so... minimum cost is it's 7. you have to solve for x, and you'll get x = 1/5(35 - i sqrt(65) ) and x = 1/5 (35 + i sqrt(65) ) which is equal to 7 -1.6 and 7 + 1.6, which is basically just 7.. ................Hope it helps.........!! :)

OpenStudy (anonymous):

Thank you so much for your help

OpenStudy (anonymous):

yeah pleasure is all mine. Anytime :)